OpenStudy (anonymous):

hey could someone please help me.... If you separate all the clubs from a deck of cards and pick 3 of those 13 cards, what is the probability that the cards are jack, queen, and king in ascending or descending order?

OpenStudy (dangerousjesse):

Firstly, do you know how many cards are in a deck, and how many different suits there are?

OpenStudy (anonymous):

4 suits 52 cards

OpenStudy (dangerousjesse):

And how many face cards?

OpenStudy (anonymous):

12 in total but 3 in clubs

OpenStudy (dangerousjesse):

Yep :) so, you have a 2/13 chance of it being either King or Jack, a 1/12 chance of it being a Queen after you choose that, and a 1/11 chance of it being either the King or the Jack

OpenStudy (anonymous):

yea..

OpenStudy (dangerousjesse):

\[\frac{2}{13}\times \frac{1}{12}\times \frac{1}{11}=\]

OpenStudy (anonymous):

its correct :D thank you

OpenStudy (anonymous):

1/858 is the answer

OpenStudy (anonymous):

sorry computer is really laggy for some reason

OpenStudy (dangerousjesse):

No problem :)

OpenStudy (anonymous):

alright have a great night :) thanks again